OpenStudy (anonymous):

I need help differentiating this : (2/3)x*x^1/2 the answer is X^1/2 , I can't see it

OpenStudy (anonymous):

Use product formula, so you have \[2/3 (\sqrt{x}) +((2/3) x \times 1/\sqrt{x}))\]

OpenStudy (anonymous):

yeahh but how does that get to\[\sqrt{x}\]

OpenStudy (anonymous):

Your answer should be \[4/3\sqrt{x}\]

OpenStudy (anonymous):

But type your question using the equation tab, cause I could have misread your question.